OpenStudy (anonymous):

verify the identity: 1) cotx/secx=cscx-sinx 2) cscx-sinx=cosxcotx 3)(1/tanx)+(1/cotx)=tanx+cotx

OpenStudy (lalaly):

have u tried to do any of them?

OpenStudy (anonymous):

yes. i've gotten stuck on the first one a couple of times. the third one i am just verifying my answer, i'm pretty sure i messed up somewhere

OpenStudy (anonymous):

1) cot/sec becomes (1/tan)/(1/cos), right?

OpenStudy (lalaly):

yep

OpenStudy (lalaly):

\[\huge{\frac{cotx}{secx}=\frac{\frac{cosx}{sinx}}{\frac{1}{cosx}}}\]\[\huge{=\frac{\cos^2x}{sinx}}=\frac{1-\sin^2x}{sinx}\]\[\huge{=\frac{1}{sinx}-\frac{\sin^2x}{sinx}}=cosecx-sinx\]